OpenStudy (anonymous):

factor and simplify sin^4x+2sin^2xcos^2x+cos^4x

jhonyy9 (jhonyy9):

us (a+b)*2=a*2 +2ab +b*2

OpenStudy (anonymous):

= (sin^2x + cos^2x )(sin^2x + cos^2x ) = 1 * 1 = 1